# RL2Grid: Benchmarking Reinforcement Learning in Power Grid Operations

### v1

@@ Abstract
Reinforcement learning (RL) can transform power grid operations by providing adaptive and scalable controllers essential for grid decarbonization. However, existing methods struggle with the complex dynamics, aleatoric uncertainty, long-horizon goals, and hard physical constraints that occur in real-world systems. This paper presents RL2Grid, a benchmark designed in collaboration with power system operators to accelerate progress in grid control and foster RL maturity. Built on a power simulation framework developed by RTE France, RL2Grid standardizes tasks, state and action spaces, and reward structures within a unified interface for a systematic evaluation and comparison of RL approaches. Moreover, we integrate real control heuristics and safety constraints informed by the operators’ expertise to ensure RL2Grid aligns with grid operation requirements. We benchmark popular RL baselines on the grid control tasks represented within RL2Grid, establishing reference performance metrics. Our results and discussion highlight the challenges that power grids pose for RL methods, emphasizing the need for novel algorithms capable of handling real-world physical systems.
@@ 1 Introduction
Power grids play a key role in efforts to combat climate change, necessitating a rapid transition to low-carbon energy and improved robustness against climate-induced extremes. This requires power grids to operate under increasing speed, scale, and uncertainty, due in large part to evolving supply and demand profiles resulting from the integration of variable renewable energy sources and distributed devices (Li et al.,, 2023). This integration creates significant challenges for human operators of power grids (Marot et al., 2022b, ), further exacerbated by the limitations of traditional power system solvers in addressing realistic systems (Chauhan et al.,, 2023). Deep reinforcement learning (RL) offers a promising approach to reshaping power grid operations, having demonstrated impressive performance in games and simulated environments over the last decade (Mnih et al.,, 2013; Silver et al.,, 2016; Wurman et al.,, 2022). However, there remain many open challenges that impede the practical application of RL in real-world environments—such as dealing with complex dynamics and aleatoric uncertainty, learning long-horizon goals, and satisfying hard physical constraints. We argue that power grids encompass many of these challenges, which are also open research questions in RL. For these reasons, investigating realistic power grid tasks from an RL perspective could yield substantial benefits for both society and the RL research community. However, progress in relevant RL methodologies is hindered by a lack of standardized benchmarks that can help promote and monitor progress, identify bottlenecks, and develop insights to address real-world challenges.
We address this gap with RL2Grid, an RL benchmark for realistic power grid operations designed in collaboration with major transmission system operators (TSOs). RL2Grid aims to accelerate progress in grid control and foster the maturity of RL methods by capturing a diverse, standardized set of increasingly complex power grid tasks that involve dealing with the combinatorially large number of possible actions available in typical grid operations. These tasks build upon RTE’s Grid2Op (Donnot,, 2020), a well-regarded simulation framework for realistic power grid operations, and are presented within a standard Gymnasium-based interface, alongside common rewards, state spaces, and action spaces to provide a common basis for comparison. We additionally present a constrained task formalization that describes the safety specifications commonly encountered in power systems (e.g., satisfying power flows, generator limits, and line limits, as well as avoiding islanding). Our collaboration with TSOs fosters the realism of RL2Grid’s tasks, and we additionally perform an in-depth analysis of RL2Grid’s design choices by investigating (e.g.) the quality of the actions available in different task settings and how the agents learn to operate the grid. We conduct a comprehensive empirical comparison of widely adopted RL algorithms on the tasks captured within RL2Grid; in particular, we evaluate several model-free RL algorithms, as they are frequently used in the literature as baselines or building blocks for more complex approaches. [^1] RL2Grid extends CleanRL (Huang et al.,, 2022) to include flexible configurations for algorithm implementation details. We also present a heuristic module facilitating the seamless incorporation of basic grid operations (e.g., line reconnection and idle actions) into the training loop of existing algorithms, which we confirm drastically improves performance and sample efficiency across all RL algorithms.
@@
Through RL2Grid, we aim to provide a launching point to foster the maturity of RL methods within real-world environments such as power grids, notably by providing realistic tasks that encompass important open questions, and by providing a standardized basis for comparative evaluation and analysis of paths forward. We further assess the effectiveness of popular learning approaches on the RL2Grid tasks. Finally, we extensively discuss important open problems in power grids and their relationship to open challenges in RL, as well as highlighting directions for further improving the realism of the power grid simulators, which is a necessary next step to enable last-mile development and deployment of the more general methodological advances we hope RL2Grid will promote.
@@ 2 Preliminaries
Power grid tasks can be defined as a Markov decision process (MDP), modeled as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\rho,R,\gamma)$; $\mathcal{S}$ and $\mathcal{A}$ are the finite sets of states and actions, respectively, $\mathcal{P}:\mathcal{S\times\mathcal{A}\times\mathcal{S}}\to[0,1]$ is the state transition probability distribution, $\rho:\mathcal{S}\to[0,1]$ is the initial uniform state distribution, $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is a reward function, and $\gamma\in[0,1)$ is the discount factor. In policy optimization algorithms, agents learn a parameterized stochastic policy $\pi:\mathcal{S}\times\mathcal{A}\to[0,1]$, modeling the probability of taking an action $a_{t}\in\mathcal{A}$ in a state $s_{t}\in\mathcal{S}$ at a certain step $t$. We can also design value-based algorithms by defining state and action value functions $V_{\pi}$ and $Q_{\pi}$ (Equation 1), which model the expected discounted return when starting from a state $s$ (and action $a$ for $Q_{\pi}$) and following the policy $\pi$ thereafter. In these contexts, agents typically use a greedy policy over the action value function (i.e., they take the action corresponding to argmax over the values). The goal is to find a policy that maximizes the expected discounted return. [^2] For the sake of clarity and brevity, we refer to RTE France, (2025) for an exhaustive overview of the MDP formalization, the state and action spaces, reward, and value ranges mentioned in the next sections.
$V_{\pi}(s)=\mathbb{E}_{\pi}\Bigg{[}\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})|s_{0}=s\Bigg{]},~{}\hfill Q_{\pi}(s,a)=\mathbb{E}_{\pi}\Bigg{[}\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})|s_{0}=s,a_{0}=a\Bigg{]}.$ (1)
@@
Constrained reinforcement learning. To enforce safety, we extend the above definition to constrained MDPs (CMDPs) (Altman,, 1998)—the most common formalism used in safe RL today (Gu et al.,, 2024). CMDPs add a set of constraints $\mathcal{C}:=\{C_{i}\}_{i=1,\ldots,n}$ that are individually defined over pairs of unsafe states and actions, identified by an indicator cost function where a positive value indicates an unsafe pair. These can be used to describe both instantaneous constraints (which must be satisfied at every point in time) and cumulative constraints (specifying limits on the total accumulation of cost over a specified horizon). While instantaneous constraints better describe the safety specifications commonly encountered in power systems (e.g., satisfying power flows and generator limits, and avoiding islanding), cumulative constraints are often easier to accommodate within existing RL optimization schemes (Su et al.,, 2024). For instance, policy optimization approaches typically transform the CMDP into an equivalent unconstrained Lagrangian optimization problem $\mathcal{L}$ over the policy parameters using dual variables as:
$\mathcal{L}^{\pi}(\bm{\lambda})=J_{R}+\mathcal{L}_{\mathcal{C}}(\bm{\lambda}),\quad\text{where}\quad\mathcal{L}_{\mathcal{C}}(\bm{\lambda})=-\sum_{i=1}^{n}\lambda_{i}(V^{\pi}_{C_{i}}-\tau_{i}),$ (2)
where the RL algorithm maximizes the return objective $J_{R}$, $\bm{\lambda}=\{\lambda_{i}\}_{i=1,\dots,n}$ act as penalties on $J_{R}$ for each constraint, $\bm{\tau}=\{\tau_{i}\}_{i=1,\dots,n}$ are the constraint thresholds, and $V^{\pi}_{\mathcal{C}}=\{V_{C_{i}}^{\pi}\}_{i=1,\dots,n}$ are the expected cost returns (defined as in Equation 1, but over cost functions). However, leveraging this formalization for power grids requires formalizing operational constraints (see Section 3.2).
2.1 Grid2Op
Grid2Op is an open-source power simulation framework designed by RTE France (France’s transmission system operator) to model power grid controllers (Donnot,, 2020). It includes realistic non-linear dynamics, uncertainty deriving from time-varying renewable energy sources, and the massive amount (combinatorially large number) of grid configurations that exist even in moderately-sized grids. Grid2Op also simulates typical operations, requiring that grids work for long horizons in a way that is robust to contingency events (e.g., disconnections, maintenance, and stochastic or “adversarial” overloads caused by extreme weather conditions), as well as adhering to physical and operational constraints. For the latter, Grid2Op models: (i) cooldown periods to prevent immediate reconnection of disconnected lines, and limits on the frequency of actions on the same line to avoid asset degradation; (ii) limited thermal capacity of transmission lines; (iii) ramp rates on generators that restrict how much power generation can change between time periods; and (iv) adherence to AC power flow constraints. Moreover, using packages like chronix2grid (Marot et al., 2020b, ), Grid2Op provides realistic time series data for load demands and generator outputs under ideal conditions.
Grid2Op has been primarily used as the base for the “Learning to Run a Power Network” (L2RPN) competitions (Marot et al., 2020a, ; Marot et al.,, 2021; Marot et al., 2022a, ). While RL methods have been used for L2RPN, they fail to provide a common ground to foster advancements in the field—each method uses custom input features and action spaces of (very) limited size, often without providing sufficient evidence on how and why these spaces were considered. Hence, to date, there is no standardized solution that allows RL researchers to easily get started in this field and compare over an established benchmark. To address this gap, our work builds on Grid2Op to provide a benchmark with standardized tasks, state and action spaces, rewards, and a safe (constrained) formalization, as well as comprehensive evaluation of common baselines inspired by L2RPN. These are critical to provide a common basis for assessing advances in RL methods (Papoudakis et al.,, 2021) as well as to improve accessibility to RL practitioners who may have limited prior knowledge of power systems. [^3] Suppl. A further discusses the relationship between L2RPN and our RL2Grid benchmark.
Figure 1: Top: Example of a discrete topological action to address an overloaded line. Bottom: Example of a continuous redispatching action to address an overloaded line.
@@ 3 RL2Grid Benchmark
RL2Grid considers the general setting of operating a power grid via topology optimization, as well as redispatch and curtailment actions, in order to keep the grid operational over a long horizon—a month of operations divided into 5 minute steps. To clarify these settings, Figure 1 shows a simplified power grid scenario consisting of four substations interconnected by transmission lines (edges), with two power generators and two loads connected to buses within each substation. Generators produce power to meet the demands (loads); the power flows through transmission lines, which also leads to power losses due to resistive heat on the lines; and substations (which may contain multiple buses) can act as “switches” to direct power flows to an extent. All of these electrical components have physical limitations: for instance, generators have ramping limits that prevent arbitrary instantaneous changes in power output, and transmission lines have maximum carrying capacities, with prolonged overloads potentially causing permanent damage and disconnections. RL2Grid allows us to address such disruptions by considering the two following categories of actions:
  • Topology optimization (Figure 1 top) involves identifying substations where a bus-split action—the type of topological action we consider—can mitigate the overload by adjusting the grid topology (i.e., how elements are currently interconnected in the grid). This approach is cost-effective for grid operators as it typically involves simple switch activation. [^4] There is some uncertainty (and debate) regarding how frequently each component can be switched safely in practice, without degrading the underlying equipment. However, determining the “optimal” topology from the combinatorial number of possible configurations is typically infeasible using existing optimization-based solvers.
  • Redispatch or curtailment (Figure 1 bottom) deals with adjusting the power flow by redispatching or curtailing the power output of fossil and renewable power generators (respectively). However, this method is often economically demanding, as it disrupts the normal operations of third parties controlling the generators and can lead to additional power costs.
In the following, we discuss the main features of RL2Grid’s grid environments (which we wrap within a traditional Gymnasium interface), and our approach to design the action and state spaces and reward function to create a standardized set of tasks. We further present a heuristic module that enables basic grid operations to be incorporated within the training loop of existing RL algorithms, and introduce the operational constraints considered in the safety-focused variants of the tasks.
@@
RL2Grid Tasks. RL2Grid tasks are designed on top of 7 main “base grids” from Grid2Op. Each of these grids has a double bus system—every electrical component (i.e., generator, load, and transmission line) has two possible connections within a substation. Table 1 summarizes the “base grids,” including their features and the number of components. These scenarios have various types of contingencies such as (i) Maintenance (M): Scheduled events included in the state. During maintenance, a line is disconnected and cannot be reconnected until maintenance is complete. (ii) Opponent (O): Unforeseen events (e.g., weather conditions) that cause a random line to disconnect. The agent does not know about these events in advance and must react in real-time (with the actions described in the following section). Once a line is disconnected, it enters a “cooldown” state, during which it cannot be reconnected for several steps. Environments may also include storage units (Batteries (B)) that can act as both generators (discharge) and loads (charge).
@@
Action spaces. Each grid has the following two types of tasks, resulting in a total of 39 tasks.
(i) Topology: Agents can take discrete actions related to topological changes (i.e., disconnecting or reconnecting a line, or changing the bus to which a component is connected). These actions are virtually free (from an economic cost perspective) since they only involve remotely activating a switch. However, they represent a significant challenge for grid operations, as the number of discrete actions scales exponentially with the number of elements connected to the substation. [^5] Human operators currently modify the grid topology manually based on historical behaviors; there is no tractable approach to obtain optimal topology optimization solutions (at scale) as of yet.
Figure 3: Ranking for the bus14 task. The dashed line separates the difficulty levels (i.e., with 50 and 209 actions).
As analyzed by Chauhan et al., (2023), the topological re-configurations for a double bus substation comprising of $N_{\text{lines}}$ lines, $N_{g}$ generators, $N_{l}$ loads, are $N=2^{N_{\text{lines}}+N_{g}+N_{l}-1}-2^{N_{g}+N_{l}}-1.$ For instance, substation #5 in Figure 2 has 2 generators, 1 load, and 4 lines (7 elements), resulting in 55 possible actions. In more complex tasks like bus36 and bus118, a single substation can have over 65,000 possible configurations. Given the large action space, we create different “difficulty levels,” in which an increasing number of topology actions are available to the agent. We selected the action spaces for these levels through extensive simulations (72 hours on the computer cluster detailed in Section 4) in which we ranked the full action space based on the survival rate for the grid. This rate represents the number of steps each action maintains the grid in normal conditions over an episode, and is specifically defined as the normalized number of steps for which an action does not cause a grid collapse (i.e., because the total demand is not satisfied or parts of the grid become disconnected). We uniformly sampled topological actions, and after ordering them by highest survival rate, we took the first $N_{\text{actions}}$ from the ordered space, where $N_{\text{actions}}$ increases with each difficulty level. [^6] To motivate our ranking method, we visually analyze the impact of the resultant action spaces in Suppl. C.1. For example, Figure 3 shows the action ordering for the bus14 grid, highlighting how all the topological actions are suitable to address grid contingencies (and how easier levels contain actions that are more likely to address a contingency). Considering these levels, RL2Grid has a total of 32 topology-based environments. [^7] Suppl. C summarizes the difficulty levels and total number of actions.
(ii) Redispatching and curtailment: Agents can take continuous actions related to changing how power generation is scheduled. Unlike topology actions, these actions are not free; economic costs arise from altering the planned generation schedule of power plants, increased fuel costs, and financial compensation for renewable energy producers, to name a few examples. Redispatching (i.e., rescheduling) actions apply to fossil fuel-based generators, while curtailment actions apply to renewable energy-based generators. Batteries, if present, are also considered generators and come with continuous actions for charging/discharging operations. This action space is relatively tractable for RL algorithms since it involves one continuous action per generator (i.e., $N=N_{g}$). Thus, we present a total of 7 continuous action-based training environments (one per base environment, in which all possible redispatching and curtailment actions are available to the agent).

### v2

@@ Abstract
Reinforcement learning (RL) can provide adaptive and scalable controllers essential for power grid decarbonization. However, RL methods struggle with power grids’ complex dynamics, long-horizon goals, and hard physical constraints. For these reasons, we present RL2Grid, a benchmark designed in collaboration with power system operators to accelerate progress in grid control and foster RL maturity. Built on RTE France’s power simulation framework, RL2Grid standardizes tasks, state and action spaces, and reward structures for a systematic evaluation and comparison of RL algorithms. Moreover, we integrate operational heuristics and design safety constraints based on human expertise to ensure alignment with physical requirements. By establishing reference performance metrics for classic RL baselines on RL2Grid’s tasks, we highlight the need for novel methods capable of handling real systems and discuss future directions for RL-based grid control. [^1] Code is available at https://github.com/emarche/RL2Grid
@@ 1 Introduction
Power grids require a rapid transition to low-carbon energy and improved robustness against climate-induced extremes in order to combat climate change. This requires operating under increasing speed, scale, and uncertainty, due in large part to evolving supply and demand profiles resulting from distributed devices and variable renewable energy sources (VREs) (Li et al.,, 2023). This integration creates significant challenges for human operators and traditional power system solvers (Marot et al., 2022b, ). To clarify what a power grid is, Figure 1 shows a simplified scenario with four substations (dots) interconnected by transmission lines (edges), two power generators, and two loads connected to buses within each substation. Generators produce power that flows through transmission lines to meet demands (loads). Transmission leads to power losses due to resistive heat on the lines, and substations (which may contain multiple buses) can act as “switches” to direct power flows to an extent. All these electrical components have physical constraints that must be satisfied (e.g., generators have ramping limits preventing arbitrary instantaneous changes in power output, and transmission lines have maximum capacities, with prolonged overloads causing disconnections and permanent damage).
Deep reinforcement learning (RL) is a promising approach for power grid operations, having demonstrated impressive control performance over the last decade (Mnih et al.,, 2013; Silver et al.,, 2016; Wurman et al.,, 2022). However, power grids encompass many open research questions in RL, including dealing with complex dynamics, aleatoric uncertainty, learning long-horizon goals, and satisfying hard physical constraints. Investigating realistic power grid tasks from an RL perspective could thus yield substantial benefits for both society and the RL research community. Nonetheless, progress in relevant RL methodologies is hindered by a lack of standardized benchmarks that can help promote and monitor progress, identify bottlenecks, and develop insights to address real-world challenges. We fill this gap by introducing RL2Grid, an RL benchmark for realistic power grid operations designed in collaboration with major transmission system operators (TSOs). RL2Grid aims to accelerate progress in grid control and advance RL methods tailored for real-world problems by modeling a diverse, standardized set of increasingly complex power grid environments for RL research. These tasks build upon RTE France’s Grid2Op (RTE France,, 2020), a realistic power grid simulation framework, and are presented within a standard Gymnasium-based interface, alongside common state and action spaces, and rewards, to provide a shared base for comparison. We additionally perform an in-depth analysis of RL2Grid’s design choices by investigating the quality of the actions available in different task settings. To incorporate power grids’ real operation practices and hard physical requirements, we also introduce a heuristic module incorporating common line reconnection and idle practices in the grid dynamics (RTE France,, 2020), as well as constrained task formalizations for safe RL (Gu et al.,, 2024). Finally, by extending the well-known CleanRL library (Huang et al.,, 2022) to include flexible configurations for algorithm implementation details, we conduct a comprehensive empirical comparison of classic RL algorithms that are frequently used in the literature as baselines or building blocks for more complex approaches. Our codebase is available as supplementary material.
Figure 1: A high-level overview of a power grid action (bus split) to address an overloaded line (red).
@@
Finally, our collaboration with TSOs allows us to extensively discuss (Section 6): (i) directions to further improve the realism of power grid simulators to enable last-mile development and deployment of the methodological advances we hope RL2Grid will promote; and (ii) the relationship of power grid operations to open challenges in RL. Through RL2Grid, we aim to foster the maturity of RL methods for real-world power grid domains and provide a standardized basis for comparative analysis.
@@ 2 Preliminaries
A power grid task can be modeled as a Markov decision process (MDP)—a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\rho,R,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are the finite sets of states and actions, respectively, $\mathcal{P}:\mathcal{S\times\mathcal{A}\times\mathcal{S}}\to[0,1]$ is the state transition probability distribution, $\rho:\mathcal{S}\to[0,1]$ is the initial uniform state distribution, $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is a reward function, and $\gamma\in[0,1)$ is the discount factor. In policy optimization algorithms, agents learn a parameterized policy $\pi:\mathcal{S}\times\mathcal{A}\to[0,1]$, modeling the probability of taking an action $a_{t}\in\mathcal{A}$ in a state $s_{t}\in\mathcal{S}$ at a certain step $t$. In value-based algorithms, agents learn state and/or action value functions $V_{\pi}$ and $Q_{\pi}$, representing the expected discounted return when starting from a state $s$ (and action $a$ for $Q_{\pi}$) and following the policy $\pi$ thereafter. In these contexts, agents typically use a greedy policy taking the action corresponding to argmax over $Q_{\pi}$. The goal is to find a policy that maximizes the expected discounted return $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$.
@@
To promote safety, we also model grid tasks as constrained MDPs (CMDPs) (Altman,, 1998), by adding a set of constraints $\mathcal{C}:=\{C_{i}\}_{i=1,\ldots,n}$ defined over unsafe state and action pairs identified by indicator cost functions. These can be used to describe both instantaneous constraints which must be satisfied at every point in time, and cumulative constraints specifying limits on the accumulation of cost over a specified horizon. For instance, policy optimization approaches typically transform the CMDP into an equivalent unconstrained Lagrangian optimization problem $\mathcal{L}$ over the policy parameters using dual variables as $\mathcal{L}^{\pi}(\bm{\lambda})=J_{R}+\mathcal{L}_{\mathcal{C}}(\bm{\lambda})$, where $\mathcal{L}_{\mathcal{C}}(\bm{\lambda})=-\sum_{i=1}^{n}\lambda_{i}(V^{\pi}_{C_{i}}-\tau_{i})$, $J_{R}$ is the return objective to maximize, $\bm{\lambda}=\{\lambda_{i}\}_{i=1,\dots,n}$ act as penalties on $J_{R}$ for each constraint, $\bm{\tau}=\{\tau_{i}\}_{i=1,\dots,n}$ are the constraint thresholds, and $V^{\pi}_{\mathcal{C}}=\{V_{C_{i}}^{\pi}\}_{i=1,\dots,n}$ are the expected cost returns.
@@ 3 RL2Grid Benchmark
RL2Grid considers the general setting of operating a power grid via topology optimization, as well as redispatch and curtailment actions (wrapped within a traditional Gymnasium interface), in order to keep the grid operational over a long horizon—a month of operations divided into 5 minute steps:
(i) Topology optimization involves identifying substations where a bus-split action can mitigate the overload by adjusting the grid topology (i.e., how elements are interconnected in the grid). This approach is cost-effective for grid operators as it typically involves simple switch activation. [^2] There is some uncertainty (and debate) regarding how frequently each component can be switched safely in practice, without degrading the underlying equipment. However, determining the “optimal” topology from the exponential number of possible configurations is typically infeasible using existing optimization-based solvers.
(ii) Redispatch or curtailment deals with adjusting the power flow by redispatching or curtailing the power output of fossil and renewable power generators (respectively). However, this method is often economically demanding, as it disrupts the normal operations of third parties controlling the generators and can lead to additional power costs.
@@
RL2Grid tasks are designed on top of 7 main “base grids” from Grid2Op. Each of these grids has a double bus system—every electrical component (i.e., generator, load, and transmission line) has two possible connections within a substation. Table 1 summarizes these base grids, along with the features and the number of components they include. These grids present two possible types of contingencies: (i) Maintenance (M): Scheduled maintenance events observed by the agent where a line is disconnected and cannot be reconnected until a fixed number of steps (a cooldown) has passed; and (ii) Opponent (O): Unforeseen events (e.g., weather conditions) that introduce stochasticity by causing a random line to disconnect, entering its cooldown state. The agent does not know about these events in advance and must address the contingencies that such a disconnection might cause in real time. Environments may also include storage units (Batteries (B)) that can act as both generators (discharge) and loads (charge).
@@
Transition dynamics. Each scenario in RL2Grid is defined over time series of synthetic but operationally realistic load demand and generation profiles created with ChroniX2Grid (Marot et al., 2020b, ). [^3] We use the time series integrated in Grid2Op’s base grids, which consists of months to years’ worth of data simulating generation and demand profiles happening over extended periods of time. Due to confidentiality and privacy constraints, up-to-date real-world grid time series data is typically unavailable to use. At the beginning of each episode, a random time index is sampled from the full-time series to initialize the environment, ensuring that agents are trained under a wide variety of grid conditions and do not overfit to specific temporal patterns.
From there, at each step, the environment transitions from state $s_{t}$ to $s_{t+1}$ through a multi-stage process that reflects realistic grid operations. First, stochastic (unforeseen) opponent events (e.g., line faults caused by weather events) are applied, following distributions specific to each base grid. Agent actions (e.g., topology or redispatch changes) are then executed, and invalid actions (e.g., those violating cooldowns) do not have any effect. The environment updates cooldown timers and applies any scheduled maintenance events. The underlying Grid2Op power simulation framework then simulates physical power flows using an AC power flow solver; if the system is infeasible due to islanding or demand shortfall, the episode ends. Otherwise, line overloads are identified, and persistent overloads lasting more than three steps trigger automated disconnections. More details on this are discussed in RTE France, (2020). The next state is constructed by updating grid variables (e.g., topology, power flows, forecasts), capturing the nonlinear, nonconvex, and stochastic dynamics of real-world power systems.